Find the solution to the following system of linear equations:
3x + 4y =-8

Not sure how to work this problem out thank you.

Found 2 solutions by Fahad_Algebra, Alan3354:
Answer by Fahad_Algebra(1)

3x+4y=-8
3x+y=-8/4
3x+y=-2
x+y=-2/3

Answer by Alan3354(69443)

Find the solution to the following system of linear equations:
3x + 4y =-8
---------------
There's only one eqn in 2 variables. It can be graphed and rearranged, but that's all.
There's nothing else that can be done with it.